# Towards Human-AI Synergy in UI Design: Enhancing Multi-Agent Based UI Generation with Intent Clarification and Alignment

### v1

@@ 4.3.3. Theme Image Generation
The objective of generating a theme image is to visually guide the overall design and ensure intent alignment. allowing the sub-agent to generate coherent and consistent design.
@@
The Text Encoder encodes the design prompt, Denoising Module denoises the original random images in several rounds, and the Autoencoder Decoder finally generates the image. ControlNet controls the generation by manipulating the denoising module by inserting an additional spatial condition to UNet. We employ UI layout as the spatial condition.
Specifically, as seen in Fig. 5, this theme image generation module takes three inputs: (a) design prompt (the text condition);(b) a latent image generated from Gaussian noise $\sim N(0,1)$ (the original image), and (c) UI layout (the spatial condition). We first put (a) design prompt into the text encoder of diffusion model, pre-trained CLIP ViT-L/14 (Frans et al., 2022) text encoder, and feed (c) UI layout to ControlNet, to obtain their embeddings respectively. We then gradually denoise (b) the latent image through the Denoising Module with the control from ControlNet. Finally, we obtain the output by feeding the denoised image into the Autoencoder Decoder.
@@
4.3.4. Sub-agent Execution
During the sub-agent execution phase, the Theme Design Agent identifies the optimal sub-agent corresponding to the component type. We considered 13 component types, as detailed by the Rico dataset. To elaborate, the Text Agent handles “Text Button” and “Text” components. The Image Agent is entrusted with “Image” and “Background Image” components, and the Icon Agent focuses on “Icon” components. For other component types, we seek to render them editable, drawing insights from RaWi (Kolthoff et al., 2023b). The colour for any component is determined by identifying the dominant RGB colour from the image region’s histogram and then representing it in HTML code.
The dynamic between the central agent and the sub-agents is essential to our system’s functionality. When a central agent engages a sub-agent, it uses a combination of past results and a cache pool to inform the sub-agents prompts:
@@ 4.3.4. Sub-agent Execution
In this context, $Res_{t-1}$ is the output from the sub-agent for the $t-1^{th}$ component. The $Cache_{t}$ represents a cache pool that integrates the previous result with accumulated knowledge from earlier iterations. This cache pool serves as an essential memory function, retaining the design context and facilitating multi-round interactions by allowing the system to “remember” past interactions and decisions. Meanwhile, $p_{t+1}$ functions as the prompt for the $t-1^{th}$ component’s sub-agent interaction, incorporating both the specific prompt $p_{sub}$ for the current sub-agent and the cumulative knowledge in $Cache_{t}$. This mechanism ensures that each sub-agent’s action is informed by the historical context, enabling consistent design, intent alignment and coherent multi-conversation interaction.
@@
Further details on how these prompts are formulated for specific sub-agents, such as the Text Content Agent and Icon Content Agent, are outlined in Sections 4.4 and 4.6, respectively.
4.4. Text Content Agent
@@
The primary role of Text Content Agent is to generate textual information tailored to specific GUI components. We use GPT-4 (OpenAi, 2024b). The system prompt for this agent, represented as $p_{text}$, is: “Based on the theme description and relevant details, provide a text content recommendation for the designated position at [bbox].” In alignment with equation 2, the execution prompt of the Text Content Agent obtains its value from the central agent’s cache, denoted as $Cache_{t-1}$, and is subsequently concatenated with $p_{text}$, to ensure consistency in the system. This systematic approach ensures that the system consistently produces text content that seamlessly integrates with the overall theme of the GUI component.
4.5. Image Content Agent
@@
To enhance the generation quality of local image-associated components and maintain the consistency of the generated outcomes, we also deploy our adaptively fine-tuned stable diffusion model in the Image Content Agent $A_{img}$. We reuse the Stable Diffusion model finetuned in Section 4.3.3 but disable the ControlNet module. Rather than using the latent image generated from Gaussian noise, we extract the area of the image component from the theme image as the input (b). In addition, we use the image description from the generated theme description as the design prompt (a). By harmonizing both textual and visual signals, we can guarantee that the produced content aligns seamlessly with the primary theme design intent.
4.6. Icon Agent
@@
The Icon Agent is crucial for selecting appropriate icons and integrating them into the graphical user interface components. In addition to acting as intuitive visual cues, well-designed icons can improve comprehension and the overall user experience. The system prompt for the Icon Agent, $p_{icon}$, is: “In reference to relevant information and taking into account its positioning at [bbox], and based on the theme description, propose an indicative phrase like “msg” for the “Icon”. As shown in equation 2, the Icon Agent’s execution prompt is based on the central agent’s cache, $Cache_{t-1}$, combined with $p_{icon}$. This approach ensures the icons selected match the GUI design semantically and visually. The Icon Agent then retrieves the optimal icon SVG code from the knowledge base, corresponding to the generated semantic phrase.
Figure 6. Interactive Interface of the PrototypeAgent Online Prototyping Editor
@@ 5. Implementation
In the implementation of the interface, we developed a web-based rapid prototyping editor using HTML5 and JavaScript, to offer a live preview of components and an interactive editing environment. Motivated by the findings from Section 3, PrototypeAgent is adept at processing design prompts and wireframe inputs to respond to F2. The Online Editor is designed for easy input, with an editable text box that let designers create task prompts seamlessly.
Adding and configuring UI elements is made simple: designers select an element type and can then resize or reposition it using a convenient drag-and-drop interface, as shown in Figure 6. The editor is structured into distinct parts for optimal user experience: Part A allows for the input of design descriptions, Part B serves as the user selection and edit panel where components such as “Web View” and “Button” can be chosen, and layout actions like “Create Layout” or “Load Layout” are available. Manual adjustments to layout size and positioning are done in Section C, while Section D provides a real-time preview of both the layout and the generated results.
@@
Beyond Section D’s preview feature, every part of our system is designed to be editable, empowering designers to efficiently refine and produce their final prototype. Upon generating components via our multi-agent system, designers can refine both the intermediate and final outputs. The system allows for real-time visualization of the application design, and any changes made by the designer are instantly updated in the system with the “Update Edit Result” button, which simplifies group text modifications and streamlines the design refinement process.
For the LLM model’s deployment, we utilized OpenAI’s gpt-4 (OpenAi, 2024b) APIs for textual generation within our multi-agent system and paired it with the text-embedding-ada-002 (OpenAi, 2024a) API for text embedding. All temperature parameters were set to 0.
Regarding the image generation model, we used the runwayml stable-diffusion-1-5 (AI, 2022) checkpoint from HuggingFace to finetune the model, and the Rico dataset. The dataset was divided into training, validation, and testing sets, consisting of 157,123 UIs, 2,364 UIs, and 4,310 UIs, respectively, following Rico (Deka et al., 2017). Training images were adjusted to $512\times 512$ resolutions, and optimization was achieved using the AdamW algorithm with a learning rate set at 1e-5 and a batch size of 1. This training utilized an Nvidia 3090 GPU 24GB VRAM.
@@ 6. Evaluation
  • RQ2: How do PrototypeAgent’s individual modules influence its performance in quality and diversity?
@@ 6.2. Metrics
(3) $FID=||\mu_{r}-\mu_{g}||_{2}^{2}+Tr(\Sigma_{r}+\Sigma_{g}-2(\Sigma_{r}\Sigma_{g})^{1/2})$
Here, $\mu_{r}$ and $\mu_{g}$ denote the mean values of the 2048-dimensional activations of the Inception-v3 pool3 layer for real ($r$) and generated samples ($g$), respectively. Meanwhile, $\Sigma_{r}$ and $\Sigma_{g}$ represent their respective covariances. A lower FID score suggests that the generated images’ distribution more closely matches that of the real images, indicating superior quality and diversity. To compute the FID score, an equal number of real and generated images are fed into the Inception-v3 (Inception, 2015) network.
Generation Diversity (GD) (Cao et al., 2019) can evaluate the diversity of generated prototypes. It calculates the pairwise distances between different UI designs within a generated set. Utilizing Perceptual Hashing (Zauner, 2010), the metric computes these distances, with larger average distances indicating a broader variety of designs within the generated set. The formula of GD is:
@@
6.3. RQ1: Comparisons to existing image generation models
6.3.1. Baselines
We consider two state-of-the-art image generation models: stable-diffusion-1-5 (AI, 2022) and stable-diffusion-2-1 (AI, 2023). Stable-diffusion-1-5, released in October 2022, is a widely accepted and stable version of the model. The subsequent version, 2-1, was unveiled in December 2022. It enhances the generation of images with greater diversity and realism, particularly for people, designs, and wildlife. Furthermore, it offers support for non-standard resolution generation. As these two baselines do not incorporate ControlNet module, we consider variants of both by integrating ControlNet, denoted as stable-diffusion-1-5 (with ControlNet), stable-diffusion-2-1 (with ControlNet). In addition, we also employ an ablated version of our approach, PrototypeAgent (w/o ControlNet) as a baseline.
6.3.2. Results
As seen in Table 4, several noteworthy observations can be made:
Dominance in Quality. Our model, PrototypeAgent, consistently outperforms both baseline models in terms of FID scores, regardless of whether ControlNet is utilized. Specifically, a lower FID score suggests that the distribution of generated images more closely matches that of real images. This indicates that PrototypeAgent’s outputs are more realistic, evident from its significantly reduced FID scores: 33.08 without ControlNet and an even lower 23.76 with ControlNet.
Consistent Diversity with Details. GD scores reflect the model’s ability to produce varied yet detailed UI designs. A higher GD indicates more detail, suggesting that the designs are diverse and intricate in their presentation. Without ControlNet, PrototypeAgent achieves a GD of 15.95, showcasing a balanced performance between quality and detailed variety. When integrated with ControlNet, the model achieves a commendable GD of 13.98. Although slightly reduced, this score underscores PrototypeAgent’s ability to produce diverse and detailed outputs, even within layout constraints. Notably, PrototypeAgent’s GD scores, both with and without ControlNet, surpass those of the baseline models. This increase in GD values emphasizes our model’s superior capability in producing designs that are varied and enriched with details compared to its peers.
Impact of ControlNet. Incorporating ControlNet results in noticeable improvements in FID scores for all models. For PrototypeAgent, the FID shows an enhancement of 10.68, representing a substantial 32% improvement. However, the slight decrease in GD, from 15.95 to 13.98, suggests that while ControlNet enhances image realism, it might limit the detail in generative diversity. This trade-off between quality and diversity is anticipated since layout constraints naturally reduce the range of potential outputs.
Comparing Our Fine-tuned Model with Baselines. Contrasted with the baseline models, the advantages of our fine-tuned model become clear. In both scenarios, with and without ControlNet, PrototypeAgent achieves superior FID scores, highlighting its excellence in UI design generation. The consistent GD scores, even surpassing some baselines, confirm the model’s capacity to generate designs that are of high quality and rich in detail. We will provide qualitative evaluation through a user study in Section 7.1.
Answer to RQ1: Our PrototypeAgent outperforms the baseline models in terms of both quality and diversity. The addition of ControlNet optimizes this performance further by introducing layout constraints, reinforcing its potential for generating realistic and detail-oriented UI designs.
6.4. RQ2: Ablation Study
6.4.1. Baselines and Ablation Strategy
To better understand the interaction and individual impact of the decoupled generation mechanisms of PrototypeAgent, we perform an ablation study, sequentially removing each module and evaluating the effect. As seen in Table 4, we carefully crafted six ablations. These modules span across the four systematic phases, namely Knowledge Retrieval, Theme Description Generation, Theme Image Generation, and Sub-agent Execution. The Sub-agent Execution phase, being more granular, was further subdivided to examine the effects of the Text Content Agent, Image Content Agent, and Icon Content Agent individually.
6.4.2. Results
Figure 7. Generated GUI Examples from Input Layout and Prompt: Corresponding Rico Screenshots, stable-diffusion-1-5 (with ControlNet), stable-diffusion-2-1 (with ControlNet), and Our PrototypeAgent.
Knowledge Retrieval. Fundamental to PrototypeAgent, this module imports essential data from our knowledge base. Excluding it results in an increase of 18.8 in the FID score (from 23.76 to 42.56). This represents the most significant decline in generation quality, underscoring that our knowledge base is pivotal in laying the foundation for high-quality design outputs.
Sub-agent Execution. This phase ensures that suitable sub-agents are selected for collaboration, optimizing each design component’s realization. Based on the evaluation of the Rico dataset, these sub-agents have specialized roles: the Text Agent managed “Text Button” and ”Text” components; the Image Agent was responsible for ”Image” and ”Background Image”; while the Icon Agent was tailored for ”Icon” components. The slight variations in FID scores — 24.06(+0.3), 24.71(+0.95), and 24.32(+0.56) — when these agents were individually excluded, it became evident from their collective contributions that our fine-grained generation approach is key to driving the model’s peak performance.
Answer to RQ2: Every module in PrototypeAgent plays a crucial role in ensuring superior design quality and diversity. The “Knowledge Retrieval” and “Theme Image Generation” modules are the most influential modules. Meanwhile, the “Theme Description Generation” and various “Sub-agents” work in concert to fine-tune and enrich the final output, adding layers of complexity and refinement to the overall design. This experiment confirms the effectiveness of our decoupling strategy, where specialized agents handle distinct components of the process.
@@ 6.4.2. Results
PrototypeAgent 23.76 13.98
Table 4. Results of the ablation study for the different modules in PrototypeAgent.
@@
-Text Content Agent 24.06 13.78
-Image Content Agent 24.71 13.38
-Icon Content Agent 24.32 13.41

### v3

@@ 4.3.3. Theme Image Generation
The objective of generating a theme image is to visually guide the overall design and ensure intent alignment. allowing the sub-module to generate coherent and consistent design.
@@
The Text Encoder encodes the prompt, Denoising Module denoises the original random images in several rounds, and the Autoencoder Decoder finally generates the image. ControlNet controls the generation by manipulating the denoising module by inserting an additional spatial condition to UNet. We employ UI layout as the spatial condition.
Specifically, as seen in Fig. 6, this theme image generation module takes three inputs: (a) prompt (the text condition);(b) a latent image generated from Gaussian noise $\sim N(0,1)$ (the original image), and (c) UI layout (the spatial condition). We first put (a) prompt into the text encoder of diffusion model, pre-trained CLIP ViT-L/14 (Frans et al., 2022) text encoder, and feed (c) UI layout to ControlNet, to obtain their embeddings respectively. We then gradually denoise (b) the latent image through the Denoising Module with the control from ControlNet. Finally, we obtain the output by feeding the denoised image into the Autoencoder Decoder.
@@
4.3.4. Sub-module Execution
During the sub-module execution phase, the Theme Design Module identifies the optimal sub-module corresponding to the component type. We considered 13 component types, as detailed by the Rico dataset. To elaborate, the Text Module handles “Text Button” and “Text” components. The Image Module is entrusted with “Image” and “Background Image” components, and the Icon Module focuses on “Icon” components. For other component types, we seek to render them editable, drawing insights from RaWi (Kolthoff et al., 2023b). The colour for any component is determined by identifying the dominant RGB colour from the image region’s histogram and then representing it in HTML code.
The dynamic between the central module and the sub-modules is essential to our system’s functionality. When a central module engages a sub-module, it uses a combination of past results and a cache pool to inform the sub-modules prompts:
@@ 4.3.4. Sub-module Execution
In this context, $Res_{t-1}$ is the output from the sub-module for the $t-1^{th}$ component. The $Cache_{t}$ represents a cache pool that integrates the previous result with accumulated knowledge from earlier iterations. This cache pool serves as an essential memory function, retaining the design context and facilitating multi-round interactions by allowing the system to “remember” past interactions and decisions. Meanwhile, $p_{t+1}$ functions as the prompt for the $t-1^{th}$ component’s sub-module interaction, incorporating both the specific prompt $p_{sub}$ for the current sub-module and the cumulative knowledge in $Cache_{t}$. This mechanism ensures that each sub-module’s action is informed by the historical context, enabling consistent design, intent alignment and coherent multi-conversation interaction. The caching mechanism is maintained and accessible by the Central Module only, who is responsible for triggering the corresponding module by providing relevant information and receiving data from each module and storing it in the cache pool.
@@
Further details on how these prompts are formulated for specific sub-modules, such as the Text Content Module and Icon Content Module, are outlined in Sections 4.4 and 4.6, respectively.
4.4. Text Content Module
@@
The primary role of Text Content Module is to generate textual information tailored to specific GUI components. We use GPT-4 (OpenAi, 2024b). The system prompt for this module, represented as $p_{text}$, is: “Based on the theme description and relevant details, provide a text content recommendation for the designated position at [bbox].” In alignment with equation 2, the execution prompt of the Text Content Module obtains its value from the central module’s cache, denoted as $Cache_{t-1}$, and is subsequently concatenated with $p_{text}$, to ensure consistency in the system. This systematic approach ensures that the system consistently produces text content that seamlessly integrates with the overall theme of the GUI component.
4.5. Image Content Module
@@
To enhance the generation quality of local image-associated components and maintain the consistency of the generated outcomes, we also deploy our adaptively fine-tuned stable diffusion model in the Image Content Module $M_{img}$. We reuse the Stable Diffusion model finetuned in Section 4.3.3 but disable the ControlNet module. Rather than using the latent image generated from Gaussian noise, we extract the area of the image component from the theme image as the input (b). In addition, we use the image description from the generated theme description as the prompt (a). By harmonizing both textual and visual signals, we can guarantee that the produced content aligns seamlessly with the primary theme design intent.
4.6. Icon Module
@@
The Icon Module is crucial for selecting appropriate icons and integrating them into the graphical user interface components. In addition to acting as intuitive visual cues, well-designed icons can improve comprehension and the overall user experience. The system prompt for the Icon Module, $p_{icon}$, is: “In reference to relevant information and taking into account its positioning at [bbox], and based on the theme description, propose an indicative phrase like “msg” for the “Icon”. As shown in equation 2, the Icon Module’s execution prompt is based on the central module’s cache, $Cache_{t-1}$, combined with $p_{icon}$. This approach ensures the icons selected match the GUI design semantically and visually. The Icon Module then retrieves the optimal icon SVG code from the knowledge base, corresponding to the generated semantic phrase.
Figure 7. Interactive Interface of the PrototypeFlow Online Prototyping Editor
@@ 5. Implementation
In the implementation of the interface, we developed a web-based rapid prototyping editor using HTML5 and JavaScript, to offer a live preview of components and an interactive editing environment. Motivated by the findings from Section 3, PrototypeFlow is adept at processing prompts and wireframe inputs to respond to F2. The Online Editor is designed for easy input, with an editable text box that let designers create task prompts seamlessly.
Adding and configuring UI elements is made simple: designers select an element type and can then resize or reposition it using a convenient drag-and-drop interface, as shown in Fig. 7. The editor is structured into distinct parts for optimal user experience: Part A allows for the input of design descriptions, Part B serves as the user selection and edit panel where components such as “Web View” and “Button” can be chosen, and layout actions like “Create Layout” or “Load Layout” are available. Manual adjustments to layout size and positioning are done in Section C, while Section D provides a real-time preview of both the layout and the generated results.
@@
Beyond Section D’s preview feature, every part of our system is designed to be editable, empowering designers to efficiently refine and produce their final prototype. Upon generating components via our multi-module system, designers can refine both the intermediate and final outputs. The system allows for real-time visualization of the application design, and any changes made by the designer are instantly updated in the system with the “Update Edit Result” button, which simplifies group text modifications and streamlines the design refinement process.
For the LLM model’s deployment, we utilized OpenAI’s gpt-4 (OpenAi, 2024b) APIs for textual generation within our multi-module system and paired it with the text-embedding-ada-002 (OpenAi, 2024a) API for text embedding. All temperature parameters were set to 0.
Regarding the image generation model, we used the runwayml stable-diffusion-1-5 (AI, 2022) checkpoint from HuggingFace to finetune the model, and the Rico dataset. The dataset was divided into training, validation, and testing sets, consisting of 15,743 UIs, 2,364 UIs, and 4,310 UIs, respectively, following Screen2Words (Wang et al., 2021). Training images were adjusted to $512\times 512$ resolutions, and optimization was achieved using the AdamW algorithm with a learning rate set at 1e-5 and a batch size of 1. This training utilized an Nvidia 3090 GPU 24GB VRAM.
@@ 6. Evaluation
  • RQ2: How do PrototypeFlow’s individual modules influence its performance in quality and diversity?
@@ 6.2. Metrics
[…]
Here, $\mu_{r}$ and $\mu_{g}$ denote the mean values of the 2048-dimensional activations of the Inception-v3 pool3 layer for real ($r$) and generated samples ($g$), respectively. Meanwhile, $\Sigma_{r}$ and $\Sigma_{g}$ represent their respective covariances. A lower FID score suggests that the generated images’ distribution more closely matches that of the real images, indicating superior quality and diversity. To compute the FID score, an equal number of real and generated images are fed into the Inception-v3 (Inception, 2015) network. This standard evaluation protocol allows for consistent and comparable results within the image generation community.
Generation Diversity (GD) (Cao et al., 2019) measures the low-level visual diversity among generated prototypes, rather than high-level semantic differences. This metric helps ensure that outputs are not overly uniform or lacking in content. GD is particularly valuable for detecting failure cases where the generative model produces nearly blank images or outputs with minimal color variation (e.g., completely black or white images, which result in low GD values). It calculates the pairwise distances between different UI designs within a generated set. Utilizing Perceptual Hashing (Zauner, 2010), the metric computes these distances, with larger average distances indicating a broader variety of designs within the generated set. The formula of GD is:
@@
PrototypeFlow 23.76 13.98
Table 4. Results of the ablation study for the different modules in PrototypeFlow.
@@
-Text Content Module 24.06 13.78
-Image Content Module 24.71 13.38
-Icon Content Module 24.32 13.41
6.3. RQ1: Comparisons to existing image generation models
6.3.1. Baselines
We consider two state-of-the-art image generation models: stable-diffusion-1-5 (AI, 2022) and stable-diffusion-2-1 (AI, 2023). Stable-diffusion-1-5, released in October 2022, is a widely accepted and stable version of the model. The subsequent version, 2-1, was unveiled in December 2022. It enhances the generation of images with greater diversity and realism, particularly for people, designs, and wildlife. Furthermore, it offers support for non-standard resolution generation. As these two baselines do not incorporate ControlNet module, we consider variants of both by integrating ControlNet, denoted as stable-diffusion-1-5 (with ControlNet), stable-diffusion-2-1 (with ControlNet). In addition, we also employ an ablated version of our approach, PrototypeFlow (w/o ControlNet) as a baseline.
6.3.2. Results
As seen in Table 4, several noteworthy observations can be made:
Dominance in Quality. Our model, PrototypeFlow, consistently outperforms both baseline models in terms of FID scores, regardless of whether ControlNet is utilized. Specifically, a lower FID score suggests that the distribution of generated images more closely matches that of real images. This indicates that PrototypeFlow’s outputs are more realistic, evident from its significantly reduced FID scores: 33.08 without ControlNet and an even lower 23.76 with ControlNet.
Consistent Diversity with Details. GD scores reflect the model’s ability to produce varied yet detailed UI designs. A higher GD indicates more detail, suggesting that the designs are diverse and intricate in their presentation. Without ControlNet, PrototypeFlow achieves a GD of 15.95, showcasing a balanced performance between quality and detailed variety. When integrated with ControlNet, the model achieves a commendable GD of 13.98. Although slightly reduced, this score underscores PrototypeFlow’s ability to produce diverse and detailed outputs, even within layout constraints. Notably, PrototypeFlow’s GD scores, both with and without ControlNet, surpass those of the baseline models. This increase in GD values emphasizes our model’s superior capability in producing designs that are varied and enriched with details compared to its peers.
Impact of ControlNet. Incorporating ControlNet results in noticeable improvements in FID scores for all models. For PrototypeFlow, the FID shows an enhancement of 10.68, representing a substantial 32% improvement. However, the slight decrease in GD, from 15.95 to 13.98, suggests that while ControlNet enhances image realism, it might limit the detail in generative diversity. This trade-off between quality and diversity is anticipated since layout constraints naturally reduce the range of potential outputs.
Comparing Our Fine-tuned Model with Baselines. Contrasted with the baseline models, the advantages of our fine-tuned model become clear. In both scenarios, with and without ControlNet, PrototypeFlow achieves superior FID scores, highlighting its excellence in UI design generation. The consistent GD scores, even surpassing some baselines, confirm the model’s capacity to generate designs that are of high quality and rich in detail. We will provide qualitative evaluation through a user study in Section 7.1.
Answer to RQ1: Our PrototypeFlow outperforms the baseline models in terms of both quality and diversity. The addition of ControlNet optimizes this performance further by introducing layout constraints, reinforcing its potential for generating realistic and detail-oriented UI designs.
6.4. RQ2: Ablation Study
6.4.1. Baselines and Ablation Strategy
To better understand the interaction and individual impact of the decoupled generation mechanisms of PrototypeFlow, we perform an ablation study, sequentially removing each module and evaluating the effect. As seen in Table 4, we carefully crafted six ablations. These modules span across the four systematic phases, namely Knowledge Retrieval, Theme Description Generation, Theme Image Generation, and Sub-module Execution. The Sub-module Execution phase, being more granular, was further subdivided to examine the effects of the Text Content Module, Image Content Module, and Icon Content Module individually.
6.4.2. Results
Figure 8. Ablation Study: GUI examples generated from input layout and prompt, with corresponding Rico screenshots, outputs from our PrototypeFlow, and results from ablated modules (Retrieved Knowledge Items, Theme/Content Modules). Green boxes and numbered circles indicate missing components; yellow indicate components with theme color inconsistencies, both due to module removal.
Knowledge Retrieval. Fundamental to PrototypeFlow, this module imports essential data from our knowledge base. Excluding it results in an increase of 18.8 in the FID score (from 23.76 to 42.56). This represents the most significant decline in generation quality, underscoring that our knowledge base is pivotal in laying the foundation for high-quality design outputs.
Fig. 8 shows that removing Retrieved Knowledge Items increases randomness in theme color and component selection. This increased diversity is highly valuable during early ideation and when creating designs from scratch, as it provides designers with a broad range of options. However, it conflicts with designers’ expectations once themes and functional requirements have been established, because the output no longer reflects the existing design knowledge base. In a real production scenario, this would force designers to perform additional manual editing.
When Theme Description Generation is removed in Fig. 8, the overall color palette remains correct because the Theme Image still enforces the primary theme color (red in the first example, blue in the second). However, this ablation introduces localized color disharmonies in specific components. Yellow boxes 1–3 and 6–7 highlight elements whose colors no longer align with the top bar or theme: for instance, black text appears where white is expected for annotated (1), a yellow icon instead of white for annotated (6), and a black icon background instead of blue for annotated (7).
Fig. 8 illustrates the effects of removing Theme Image Generation. We can observe that the global theme color remains consistent because the textual description still works on the palette. However, local accessibility problems emerge. For example, in yellow boxes 4 to 5, a white icon rendered on a white search bar background becomes invisible, and other elements lack the contrast needed for a further improvement. These results demonstrate that textual cues alone are insufficient: the image module is an important module for achieving those cues in a consistent, accessible, and visually compelling interface.
Sub-module Execution. This phase ensures that suitable sub-modules are selected for collaboration, optimizing each design component’s realization. Based on the evaluation of the Rico dataset, these sub-modules have specialized roles: the Text Module managed “Text Button” and ”Text” components; the Image Module was responsible for ”Image” and ”Background Image”; while the Icon Module was tailored for ”Icon” components. The slight variations in FID scores — 24.06(+0.3), 24.71(+0.95), and 24.32(+0.56) — when these modules were individually excluded, it became evident from their collective contributions that our fine-grained generation approach is key to driving the model’s peak performance.
Ablating the Text, Image, or Icon Content Modules confirms their specialized roles. Compared with full PrototypeFlow output, green boxes 1–24 in Fig. 8 mark missing text, image, or icon components when the corresponding module is disabled.
Answer to RQ2: Every module in PrototypeFlow plays a crucial role in ensuring superior design quality and diversity. The “Knowledge Retrieval” and “Theme Image Generation” modules are the most influential modules. Meanwhile, the “Theme Description Generation” and various “Sub-modules” work in concert to fine-tune and enrich the final output, adding layers of complexity and refinement to the overall design. This experiment confirms the effectiveness of our decoupling strategy, where specialized modules handle distinct components of the process.